OpenStudy (anonymous):

Evaluate sin(sin^-1(3/4)+cos^-1(1/4))without using a calculator.

OpenStudy (anonymous):

Let : \(sin^{-1}(\frac{3}{4}) = \theta\) and \( cos^{-1}(\frac{1}{4}) = x\) First solving for first part: \[\sin^{-1}(\frac{3}{4}) = \theta\] \[\sin(\theta) = \frac{3}{4}\] \[\cos(\theta) = \sqrt{1 - \sin^2(\theta)} \implies \cos(\theta) = \sqrt{1 - (\frac{3}{4})^2} \implies \cos(\theta) = \frac{\sqrt{7}}{4}\]

OpenStudy (anonymous):

Similarly can you do it for second part: \[\cos^{-1}(\frac{1}{4}) = x\] can you find sin(x) here?

OpenStudy (anonymous):

@sfb you there?